OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (3, 5)?

OpenStudy (anonymous):

y = (x – 5)2 + 3 y = (x + 5)2 + 3 y = (x – 3)2 + 5 y = (x + 3)2 + 5

OpenStudy (anonymous):

y=+-a(x-3)2 +5

OpenStudy (slaaibak):

Third one.

OpenStudy (anonymous):

Factor completely. 16x2 – 4 2(2x – 1)2 2(2x + 1)2 4(2x – 1)(2x – 1) 4(2x – 1)(2x + 1)